Suppose that the probability of Sam passing his Business Statistics exam is 0.60 and the probability that he passes his Law exam is 0.65. If passing his Business Statistics exam and passing his Law exam are independent events, then the probability (to three decimal places) that he passes his Business Statistics exam or does not pass his Law exam is?

1
Expert's answer
2021-11-16T11:27:32-0500

"P(Business\; and\; not Law)=0.60*(1-0.65)=0.21."

"P(Business\;or\;not Law)=0.60+(1-0.65)-0.21=0.74."